OpenStudy (anonymous):

Find the zeros by factoring. 1) l(x)=x^2+7x+10

jhonyy9 (jhonyy9):

you need getting two numbers with sum equal 7 and product equal 10 - ok,?

jhonyy9 (jhonyy9):

let being a and b so than you can rewriting this quadratic (x+a)(x+b) so than i give you the a you will can give the b ?

jhonyy9 (jhonyy9):

so a=2 than you know that 2+b=7 and 2*b=10 - so what will be b=?

OpenStudy (anonymous):

will it be 5?

jhonyy9 (jhonyy9):

yes right so you know that a=2 and b=5 so how will be factored this quadratic ?

jhonyy9 (jhonyy9):

how i have wrote above (x+a)(x+b) = x^2 +7x +10 --- ok. ?

jhonyy9 (jhonyy9):

hope these was helpfully - bye

OpenStudy (anonymous):

yes, it was very helpful. Thank you